# Polynomial Time Convergence of the Iterative Evaluation of Datalogo Programs

### 21 Dec 2023 v1

@@ Abstract
1 Introduction
@@
Motivated by the need in modern data analytics to express recursive computations with aggregates [10] introduced $\text{\sf Datalog}^{\circ}$, which is an extension of Datalog that allows for aggregation and recursion over an arbitrary commutative semiring. Like Datalog, $\text{\sf Datalog}^{\circ}$ programs can be evaluated via the natural iterative algorithm until a fixed point is reached. The number of iterations until the iterative evaluation of a Datalog program converges is at most the output size. In contrast the natural iterative evaluation of some $\text{\sf Datalog}^{\circ}$ programs over some commutative semirings may not converge. It is known that the commutative semirings for which the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is guaranteed to converge are exactly those semirings that are stable [10]. A semiring is $p$-stable if the number of iterations required for any one-variable recursive linear $\text{\sf Datalog}^{\circ}$ program to reach a fixed point is at most $p$, and a semiring is stable if there exists a $p$ for which it is $p$-stable. Previously, the best known upper bound on the number of iterations until convergence is $\sum_{i=1}^{n}(p+2)^{i}=\Theta(p^{n})$ steps, where $n$ is (essentially) the output size, and $p$ is stability of the underlying semiring. In contrast there are no known lower bounds that show that iterative evaluation requires an exponential (in the parameter $n$) number of steps to reach convergence.
Many naturally arising $\text{\sf Datalog}^{\circ}$ programs are linear, which are those where each product only contains at most one output database. [9] showed that iterative evaluation of all linear $\text{\sf Datalog}^{\circ}$ programs converges after $O(pn^{3})$ steps. So iterative evaluation of linear $\text{\sf Datalog}^{\circ}$ programs over stable commutative semirings converges in a polynomial number of steps. Further, [9] showed that the iterative evaluation of all linear $\text{\sf Datalog}^{\circ}$ programs over commutative $p$-stable commutative semirings containing $L$ elements converges within $O(pn\lg L)$ steps.
@@
The research question we address in this paper is whether the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs can indeed take exponential number of steps to converge. Another way to frame our motivating research question is whether or not polynomial convergence is a special property of linear $\text{\sf Datalog}^{\circ}$ programs that is not shared by general $\text{\sf Datalog}^{\circ}$ programs. Our main finding is stated in Theorem 1.1.
@@
Let $\bm{S}$ be a $p$-stable commutative semiring. Let $P$ be a $\text{\sf Datalog}^{\circ}$ program where the maximum number of multiplicands in any product is at most $\lambda$. Let $D$ be the input EDB database. Let $\sigma$ be number of the semiring elements referenced in $P$ or $D$. Let $n$ denote the total number of ground atoms in a IDB that at some point in the iterative evaluation of $P$ over $\bm{S}$ on input $D$ have a nonzero associated semiring value. Then the iterative evaluation of $P$ over $\bm{S}$ on input $D$ converges within
$pn(n+3)\cdot(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma+1)$
@@
Let us now focus on how we believe one should interpret the statement of Theorem 1.1. It is most natural to focus on how the number of iterations grows as a function of the parameters that are most likely to be large, namely the size of the input databases and the size of the output databases. $\text{\sf Datalog}^{\circ}$ programs, like Datalog programs, use the semiring zero element to signify the nonpresence of a ground tuple in a IDB. Most natural $\text{\sf Datalog}^{\circ}$ programs are monotone in the sense that a ground atom will not be deleted from an IDB after it is added (which means not resetting the associated semiring value to zero after it has become nonzero). For monotone $\text{\sf Datalog}^{\circ}$ programs, $n$ will be the number of ground atoms in the output, and we think it most natural/informative to think of $n$ in this way. Since the stability parameter $p$ is only dependent on the underlying semiring, and is independent of the input database and even in some sense independent of the $\text{\sf Datalog}^{\circ}$ program, we think it is best to view $p$ as a constant that does not scale with the input size or output size. It is natural to think of the program $P$ as being modestly sized, so it’s natural to also think of $\lambda$ as a constant that does not scale with the input size or output size. It’s natural that as the size of the input database scales that there will be significantly more semiring elements referenced in the input database than in the $\text{\sf Datalog}^{\circ}$ program. Thus it is natural to view $\sigma$ as essentially upper bounded by the number of elements in the input databases. Under these assumptions, Theorem 1.1 can be interpreted to say that the number of iterations is
$O(pn^{4}\sigma(\lg\lambda+\lg\sigma))$ (1)
So if one is in a situation where the size of the output databases are quite small compared to the size of the input databases, then the worst-case number of iterations grows at a rate at most (essentially) linearly with the size of the input databases. Otherwise, if one is in the situation where the size of the output databases are quite large compared to the size of the input database (or the input databases have a lower diversity of semiring values), then the worst case number of iterations grows proportionally to the fourth power of the size of the output databases. Finally, if one is in the situation where the size of the output databases are comparable to the size of the input databases (and the input databases have a wide diversity of semiring values), then the worst case number of iterations grows proportionally to the fifth power of the size of the output (or input) databases. Recall that [9] showed that in the worst case the number of iterations for the iterative evaluation of linear $\text{\sf Datalog}^{\circ}$ programs grows proportional to the cube of the size of the output database.
Let us assume for a moment that the number $L$ of elements in the semiring is large. Then to the extent that one can reasonably interpret $n^{4}$ (or maybe $n^{5}$) as being not too much worse than $n^{3}$, Theorem 1.1 can reasonably be interpreted as saying that the worst-case time to convergence for the iterative evaluation of general $\text{\sf Datalog}^{\circ}$ programs is not too much worse than the time to convergence for the iterative evaluation of linear $\text{\sf Datalog}^{\circ}$ programs. On the other hand, if $L$ is not too large, then it is reasonable to interpret the $O(pn\lg L)$ convergence time for linear $\text{\sf Datalog}^{\circ}$ programs in [9] to be significantly better than the bound for general $\text{\sf Datalog}^{\circ}$ programs in Theorem 1.1.
In the rest of the introduction we give a fast paced coverage of the most essential background knowledge required to understand this result ( For a deeper dive see  [10]). We then give a brief technical overview of the proof of Theorem 1.1.
1.1 Background
1.1.1 Datalog
A (traditional) Datalog program $P$ consists of a set of rules of the form:
@@ 1.1.1 Datalog
where $R_{0},\ldots,R_{m}$ are predicate names (not necessarily distinct) and each $\bm{X}_{i}$ is a tuple of variables and/or constants. The atom $R_{0}(\bm{X}_{0})$ is called the head, and the conjunction $R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{m}(\bm{X}_{m})$ is called the body. Multiple rules with the same head are interpreted as a disjunction. A predicate that occurs in the head of some rule in $P$ is called an intensional database predicate (IDB), otherwise it is called an extensional database predicate (EDB). The EDBs form the input (databases), and the IDBs represent the output (databases) computed by the Datalog program. The finite set of all constants occurring in an EDB is called the active domain, and denoted ADom. There is an implicit existential quantifier over the body for all variables that appear in the body, but not in the head, where the domain of the existential quantifier is ADom. Thus, a Datalog program can also be viewed as a collection of unions of conjunctive queries (UCQs), one UCQ for each IDB.
@@
One classic example of a Datalog program is the following one, which computes the transitive closure of a directed graph, defined by the edge relation $E(X,Y)$:
@@
Here $E$ is an EDB predicate, $T$ is an IDB predicate, and ADom consists of the vertices in the graph. Written as a UCQ, where the quantifications are explicit, this program is:
@@
A Datalog program can be thought of as a function, called the immediate consequence operator (ICO), where the ICO adds a ground atom $A$ to a IDB $R$ if there a predicate where $R$ is the head, and $R(A)$ is logically implied by the body. The iterative evaluation of a Datalog program works in rounds/steps, where on each round the ICO is applied to the current state. In the context of the Datalog program in Example 1, if $(a,b)$ was a ground atom in $T$, meaning that it known that there is a directed path from vertex $a$ to vertex $b$, and $(b,c)$ was a ground atom in $E$, meaning that there is a directed edge from $b$ to $c$, then the ICO would add the ground atom $(a,c)$ to $T$ (if it was not already present). This is because the Datalog program states that $T(a,c)$ is logically implied by $T(a,b)$ and $E(b,c)$.
1.1.2 $\text{\sf Datalog}^{\circ}$
Like Datalog programs, a $\text{\sf Datalog}^{\circ}$ program consists of a set of rules, where the union of conjunctive queries are now replaced with sum-product queries over a commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$. Each rule has the form:
@@ 1.1.2 $\text{\sf Datalog}^{\circ}$
where sum is over the active ADom of the variables not in $X_{0}$. Further each ground atom in an EDB predicate or IDB predicate is associated with an element of the semiring $\bm{S}$, and the element associated with a tuple in an EDB predicate is specified in the input. The EDBs form the input database, and the ground atoms in the IDB’s that have an associated semiring value that is nonzero represent the output instance computed by the $\text{\sf Datalog}^{\circ}$ program. Note that in a Datalog program the ground atom present in one of the databases can be thought of as those that are associated with the element $1$ in the standard Boolean semiring.
@@
One important class of a $\text{\sf Datalog}^{\circ}$ programs is those of the following form:
@@
Here $E$ is an EDB predicate, and $T$ is an IDB predicate. Note that this $\text{\sf Datalog}^{\circ}$ program is the Datalog program given in line (3) with $\vee$ replaced by $\oplus$, $\wedge$ replaced by $\otimes$ and $\exists_{Z}$ replaced by $\bigoplus_{Z}$. Here $\oplus$ is the semiring addition operation, and $\otimes$ is the semiring multiplication operation and $\bigoplus_{Z}$ is aggregation, that is an iterative application of $\oplus$ over ADom. Note that when interpreted over the Boolean semiring, this $\text{\sf Datalog}^{\circ}$ program is essentially the transitive closure program from Example 1.
@@
When interpreted over the tropical semiring $\text{\sf Trop}^{+}=(\mathbb{R}_{+}\cup\{\infty\},\min,+,\infty,0)$, the $\text{\sf Datalog}^{\circ}$ program in Example 2 solves the classic All-Pairs-Shortest-Path problem (APSP), which computes the shortest path length $T(X,Y)$ between all pairs $X,Y$ of vertices in a directed graph specified by an edge relation $E(X,Y)$, where the semiring element associated with $E(X,Y)$ is the length of the directed edge $(X,Y)$.
@@
A $\text{\sf Datalog}^{\circ}$ program can be thought of as an immediate consequence operator (ICO). To understand how the ICO works in $\text{\sf Datalog}^{\circ}$, consider a rule with with head $R$ and let $A$ be a ground atom for $R$ with associated semiring value $y$, and assume that for $A$ the body of this rule evaluates to the semiring element $x$. As a result of this, the ICO adds $x$ to the semiring value associated with $A$, resulting in a new value of $x+y$. Note that the functioning of the $\text{\sf Datalog}^{\circ}$ ICO, when the semiring is the standard Boolean semiring, is identical to how the Datalog ICO functions. The iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program works in rounds/steps, where initially the semiring element 0 is associated with each possible ground atom in an IDB, and on each round the ICO is applied to the current state. So in the context of the $\text{\sf Datalog}^{\circ}$ program in Example 3, if $(a,b)$ was a ground atom in $T$ with associated semiring element $x$, meaning that the shortest known directed path from vertex $a$ to vertex $b$ has length $x$, and $(b,c)$ was a ground atom in $E$ with associated semiring element $y$, meaning that there is a directed edge from $b$ to $c$ with length $y$, then the ICO would make the semiring element associated with $A$ the minimum (as minimum is the addition operation in the tropical semiring) of its current value and $x+y$ (as normal additional is the multiplication operation in the tropical semiring).
A $\text{\sf Datalog}^{\circ}$ program is linear if every rule has no more than one IDB prediciate in its body. The $\text{\sf Datalog}^{\circ}$ program in Example 2 is linear. While many natural $\text{\sf Datalog}^{\circ}$ programs are linear, there are also natural nonlinear $\text{\sf Datalog}^{\circ}$ programs.

### 21 Feb 2024 v2

@@ Abstract.
Key words and phrases:
Datalog, convergence time, semiring
1. Introduction
Motivated by the need in modern data analytics to express recursive computations with aggregates, Khamis et al. [7] introduced $\text{\sf Datalog}^{\circ}$, which is an extension of Datalog that allows for aggregation and recursion over an arbitrary commutative semiring. [^1] The results in [7] are on Partially Ordered Pre-Semirings (POPS). However, the key convergence properties are reflected in the core semiring of the POPS. Thus, it is sufficient to restrict our attention to semirings for the purpose of this paper. Like Datalog, $\text{\sf Datalog}^{\circ}$ programs can be evaluated via the natural iterative algorithm until a fixed point is reached. This is sometimes called the “naïve evaluation” algorithm. Furthermore, $\text{\sf Datalog}^{\circ}$ is attractive for practical applications because it also allows for a generalization of semi-naïve evaluation to work, under some assumptions about the semiring [7]. While semi-naïve evaluation makes each iteration faster to compute, the total number of iterations is the same as that of the naïve evaluation algorithm. Thus, bounding the number of iterations of the naïve evaluation is an important question in practice.
In Datalog, it is easy to see that the number of iterations until a fixed point is reached is at most the output size. (Every iteration before convergence must derive at least one new fact, due to monotonicity.) In contrast, the naïve evaluation of $\text{\sf Datalog}^{\circ}$ programs over some commutative semirings may not converge. (A simple example is the sum-product semiring over the reals.)
@@ 1. Introduction
It is known that the commutative semirings for which the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is guaranteed to converge are exactly those semirings that are stable [7]. A semiring is $p$-stable [5] if the number of iterations required for any one-variable recursive linear $\text{\sf Datalog}^{\circ}$ program to reach a fixed point is at most $p$, and a semiring is stable if there exists a $p$ for which it is $p$-stable. Previously, the best known upper bound on the number of iterations until convergence is $\sum_{i=1}^{n}(p+2)^{i}=\Theta((p+2)^{n})$ steps, where $n$ is (essentially) the output size, and $p$ is the stability index of the underlying semiring. In contrast there are no known lower bounds that show that iterative evaluation requires an exponential (in the parameter $n$) number of steps to reach convergence.
There are special cases where polynomial convergence rate is known. The first case is when the semiring is 00-stable, as in the standard Boolean semiring, where it is known that naïve evaluation converges in $O(n)$ steps [7]. The second case is when the input program is linear, meaning that in every rule the product only contains at most one IDB relational symbol. In [6] it is shown that if the semiring is $p$-stable with $L$ elements in the semiring domain, then the iterative evaluation of all linear $\text{\sf Datalog}^{\circ}$ programs converge after $O(\min(pn^{3},pn\lg L))$ steps.
@@
Our Contributions. The open problem we address in this paper is whether the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings might indeed require an exponential number of steps to converge. Another way to frame our motivating research question is whether or not polynomial convergence is a special property of linear $\text{\sf Datalog}^{\circ}$ programs that is not shared by general $\text{\sf Datalog}^{\circ}$ programs. Our main finding is stated in Theorem 1.1.
@@
Let $\bm{S}$ be a $p$-stable commutative semiring. Let $P$ be a $\text{\sf Datalog}^{\circ}$ program where the maximum number of multiplicands in any product is at most $\lambda$. Let $D$ be the input database instance. Let $\sigma$ be number of the semiring elements referenced in $P$ or $D$. Let $n$ denote the total number of ground atoms in an IDB that at some point in the iterative evaluation of $P$ over $\bm{S}$ on input $D$ have a nonzero associated semiring value. Then the iterative evaluation of $P$ over $\bm{S}$ on input $D$ converges within
$\lceil pn(n+3)\cdot(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma+1)\rceil$
@@
As $\lambda$ is only a property of the $\text{\sf Datalog}^{\circ}$ program, and $p$ is only a property of the semiring, they do not scale with the data size. Thinking of them as constants in data complexity, the bound in the Theorem 1.1 is reduced to $O(n^{4}\sigma\lg\sigma)$. Note that $\sigma$ is bounded by the input size plus the query size and so it is linear in the input size under combined complexity. [^2] Note that in [7] the parameter $\sigma$ denoted the number of references to semiring elements, not the number of semiring elements referenced. The maximum number of ground IDB atoms is $\tilde{O}(|D|^{k})$ where $k$ is the maximum arity of IDB atoms, where $\tilde{O}$ hides query-dependent factors. Thus, overall, in data complexity Theorem 1.1 gives a polynomial bound on the convergence rate; furthermore, its dependency on the output size makes the bound more flexible.
Related Works. There is a large body of research on fixed points of multi-variate polynomial functions over semirings, which were studied by many communities since the 1960s. (See e.g. [10, 7, 4, 13].) In some special cases, such as closed semirings or $\omega$-continuous semirings, there are non-iterative methods to find the fixpoint [11, 4]. In general, the iterative algorithm is still the most general.
The two papers in the literature that we directly build on are [7] and [6]. In [7] it is shown that the naïve evalution of $\text{\sf Datalog}^{\circ}$ programs converges in $O((p+2)^{n})$ steps; this bound is obtained by showing how to bound the convergence time for a high dimensional function in terms of the convergence time for a 1-dimensional function.
The paper [6] considers linear $\text{\sf Datalog}^{\circ}$ programs, where the grounded ICO is a linear function $\bm{f}:S^{n}\rightarrow S^{n}$ that can expressed as $\bm{f}(x)=A\otimes x\oplus b$ where $A$ is an $n$ by $n$ matrix with entries from the semiring. Then $f_{i}^{(q)}(0)$ becomes $\oplus_{W\in\mathcal{W}^{i}_{q}}Z(W)$, where $\mathcal{W}^{i}_{q}$ is the collection of all walks starting from ground atom $i$ with at most $q$ hops in the natural complete digraph underlying $A$, and $Z(W)$ is the product of $a_{uv}$ for every edge $uv\in W$. The crux of the $O(pn^{3})$ upper bound analysis in [6] was then that any walk $W$ longer than $O(pn^{3})$ must contain a cycle $C$ where all the edges are traversed many times. The fact that adding $Z(W)$ didn’t change the sum followed from the stability of the semiring element that is the product of the semiring values on the edges in $C$. Our analysis for the nonlinear case is more involved than the analysis for the linear cases because finding the collection of semiring values that will serve the role of the cycle $C$ in the linear case is more involved. It is interesting to note that, however, the gap between the two cases is $O(n\sigma\lg\sigma)$.
Our main theorem is proved via a strengthening of Parikh’s Theorem [12], which we believe is novel [^3] However, as there are a daunting number of different statements, proofs and extensions of Parikh’s theorem in the literature [8], it is hard to be totally confident about the novelty of our extension. and may be of independent interests.
Paper Organization. Section 2 covers background knowledge required to understand this result. Section 3 gives a brief technical overview of the proof of Theorem 1.1. Sections 4,  5, and 6 give the proof details. Finally Section 7 concludes the paper.
2. Background
2.1. Semirings
A semiring is a tuple $\bm{S}=(S,\oplus,\otimes,0,1)$ where $\oplus$ and $\otimes$ are binary operators on $S$, $(S,\oplus,0)$ is a commutative monoid (meaning $\oplus$ is commutative and associative, and 00 is the identity for $\oplus$), $(S,\otimes,1)$ is a monoid (meaning $\otimes$ is associative, and 00 is the identity for $\oplus$), $a\otimes 0=0\otimes a=0$ for every $a\in S$, and $\otimes$ distributes over $\oplus$. $\bm{S}$ is said to be commutative if $\otimes$ is commutative. Define
$u^{(p)}:=1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p},$
where $u^{i}:=u\otimes u\otimes\cdots\otimes u$ ($i$ times). An element $u\in S$ is $p$-stable if $u^{(p)}=u^{(p+1)}$, and a semiring $\bm{S}$ is $p$-stable if every element $u\in S$ is $p$-stable.
A function $f:S^{n}\to S^{n}$ is $p$-stable if $f^{(p+1)}(\bm{0})=f^{(p)}(\bm{0})$, where $\bm{0}$ is the all zero vector, and $f^{(k)}$ is the $k$-fold composition of $f$ with itself. The stability index of $f$ is the smallest $p$ such that $f$ is $p$-stable. See [5] for more background on semirings and stability.
2.2. Datalog
A (traditional) Datalog [1] program $P$ consists of a set of rules of the form:
@@ 2.2. Datalog
where $R_{0},\ldots,R_{m}$ are predicate names (not necessarily distinct) and each $\bm{X}_{i}$ is a tuple of variables and/or constants. The atom $R_{0}(\bm{X}_{0})$ is called the head, and the conjunction $R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{m}(\bm{X}_{m})$ is called the body. Multiple rules with the same head are interpreted as a disjunction. A predicate that occurs in the head of some rule in $P$ is called an intensional database predicate (IDB), otherwise it is called an extensional database predicate (EDB). The EDBs form the input, and the IDBs represent the output computed by the Datalog program. The finite set of all constants occurring in an EDB is called the active domain, and denoted ADom. An atom $R(\bm{X})$ is called a ground atom if all its arguments are constants. There is an implicit existential quantifier over the body for all variables that appear in the body, but not in the head, where the domain of the existential quantifier is ADom. Thus, a Datalog program can also be viewed as a collection of unions of conjunctive queries (UCQs), one UCQ for each IDB.
@@
A classic example of a Datalog program is the transitive closure program
@@
Here $E$ is an EDB predicate, representing the edge relation of a directed graph, $T$ is an IDB predicate, and ADom is the vertex set. Written as a UCQ, where the quantifications are explicit, this program is:
@@
The UCQ format is the right format to work with when extending Datalog programs to general semirings.
A Datalog program $P$ can be thought of as a function, called the immediate consequence operator (ICO), mapping a subset of ground IDB atoms to a subset of ground IDB atoms. (The ground EDB atoms are inputs and thus remain constants.) In particular, the ICO adds a ground (IDB) atom $R(\bm{x})$ to the output if it can be logically inferred by the input ground atoms via the rules of $P$. The iterative evaluation of a Datalog program works in rounds/steps, where on each round the ICO is applied to the current state, starting from the empty state.
2.3. Datalogo
Like Datalog programs, a $\text{\sf Datalog}^{\circ}$ program consists of a set of rules, where the UCQs are replaced by sum-sum-product queries over a commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$, where $\vee$ is replaced with $\oplus$ and $\wedge$ with $\otimes$. Specifically, in a $\text{\sf Datalog}^{\circ}program$ each rule has the form:
@@ 2.3. Datalogo
where sum is over the active ADom of the variables not in $X_{0}$. Multiple rules with the same head are combined using the $\oplus$ operation, which is the analog of combining rules using $\vee$ in Datalog.
Furthermore, each ground EDB or IDB atom is associated with an element of the semiring $\bm{S}$, and the non-zero elements associated with ground EDB atoms are specified in the input. A fixed point solution to the $\text{\sf Datalog}^{\circ}$ program associates a semiring element to ground IDB atoms. Just like in Datalog, we do not have to explicitly represent the zero-assigned ground IDB atoms: every ground atom not in the output are implicitly mapped to 00.
@@
The $\text{\sf Datalog}^{\circ}$-version of the Datalog program given in line (2) with $\vee$ replaced by $\oplus$, $\wedge$ replaced by $\otimes$ and $\exists_{Z}$ replaced by $\bigoplus_{Z}$ is
@@
Here $E$ is an EDB predicate, and $T$ is an IDB predicate, $\oplus$ is the semiring addition operation, $\otimes$ is the semiring multiplication operation and $\bigoplus_{Z}$ is aggregation, that is an iterative application of $\oplus$ over ADom.
When interpreted over the Boolean semiring, the $\text{\sf Datalog}^{\circ}$ program in (4) is the transitive closure program from Example 1.
When interpreted over the tropical semiring $\text{\sf Trop}^{+}=(\mathbb{R}_{+}\cup\{\infty\},\min,+,\infty,0)$, the $\text{\sf Datalog}^{\circ}$ program in (4) solves the classic All-Pairs-Shortest-Path (APSP) problem, which computes the shortest path length $T(X,Y)$ between all pairs $X,Y$ of vertices in a directed graph specified by an edge relation $E(X,Y)$, where the semiring element associated with $E(X,Y)$ is the length of the directed edge $(X,Y)$.
[…]
A $\text{\sf Datalog}^{\circ}$ program can be thought of as an immediate consequence operator (ICO). A simple way to understand the semantics of $\text{\sf Datalog}^{\circ}$ is to think of each body predicate $R_{i}$ in (3) as a function from the domain of $\bm{X}_{i}$ to the domain $S$ of the semiring. The functional value $R_{i}(\bm{c}_{i})$ for a particular binding $\bm{c}_{i}$ in the domain of $\bm{X}_{i}$ is the value assigned to the ground atom $R_{i}(\bm{c}_{i})$. The rule (3) is thus exactly a sum-product query  (or a functional aggregate query [2] over one semiring), and multiple rules with the same head are combined into a sum-sum-product query. The $\text{\sf Datalog}^{\circ}$ program containing these queries compute new (IDB) functions from old (IDB and EDB) functions, using the sums and products from the semiring.
The iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program works by initially assigning all IDB “functions” to be identically 00 (i.e. their ground atoms are assigned with 00). The ICO is then repeatedly applied to the current IDB state. In the context of the $\text{\sf Datalog}^{\circ}$ program in (5), initially all $T(x,z)$ are assigned with +$\infty$+ ∞ (the 00 of the tropical semiring), and the rule (5) effectively is the well-known Bellman-Ford algorithm [3].
A $\text{\sf Datalog}^{\circ}$ program is linear if every rule (3) has no more than one IDB predicate in its body. The $\text{\sf Datalog}^{\circ}$ program in Example 2 is linear. While many natural $\text{\sf Datalog}^{\circ}$ programs are linear, there are also natural nonlinear $\text{\sf Datalog}^{\circ}$ programs.
@@
As a classic example of a nonlinear $\text{\sf Datalog}^{\circ}$ program, consider the following alternate formulation of APSP, which is equivalent to (5)
@@
The convergence rate of a $\text{\sf Datalog}^{\circ}$ program is the stability index of its ICO.
2.4. Grounding the ICO
Since the final associated semiring values of the ground IDB atoms are not initially known, it is natural to think of them as (IDB) variables. Then the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$ program is a map $\bm{f}:S^{n}\rightarrow S^{n}$, where $S$ is the semiring domain, and $n$ is the number of ground IDB atoms that ever have a nonzero value at some point in the iterative evaluation of the program. For instance, in (5), there would be one variable for each pair $(x,y)$ of vertices where there is a directed path from $x$ to $y$ in the graph. So the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$ program has the following form:
$-\displaystyle X_{1}\text{ :- }$- $f_{1}(X_{1},\ldots,X_{n})$
(7) $\ldots$
$-\displaystyle X_{n}\text{ :- }$- $f_{n}(X_{1},\ldots,X_{n})$
where the $X_{i}$’s are the IDB variables, and $f_{i}$ is the component of $\bm{f}$ corresponding to the IDB variable $X_{i}$. Note that each component function $f_{i}$ is a multivariate polynomial in the IDB variables of degree at most the maximum number of factors in any product in the body of some rule (3) in the $\text{\sf Datalog}^{\circ}$ program. After $q$ iterations of the iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program, the semiring value associated with the ground atom corresponding to $X_{i}$ will be:
(8) ${f}_{i}^{(q)}(\bm{0})$